A certain battery runs a toy truck and becomes partially discharged. In the process, it performs a total of 117 J of work on its immediate surroundings. It also gives off 3 J of heat,  which the surroundings absorb. No other work or heat is exchanged with the surroundings.

  1. Provide values for and explain the sign allocated to each variable.                                                         
1
Expert's answer
2021-05-09T13:07:12-0400
"Q=-3\ J\\W=-117\ J"

"\Delta U=Q+W=-3-117=-120\ J\\"